A meat inspector has randomly selected 25 packs of 98% lean beef produced by Company A. Thesample resulted in a mean of 96% with a sample standard deviation of 1%. Assume normality.

(1) Find a 99% confidence interval for the mean leanness of lean beef produced by company A.

(2) Find a 95% prediction interval for the leanness of a new pack.
Answer #1

(1)

  \\ We\ are\ 99\%\ confident\ that\ \\ \\ 95.4\%<\mu<69.6\%

(b)

\\ 95\%\ prediction\ interval\ is:\\ \\ \left(\bar{x}-t_{\alpha}s\left(\sqrt{1+\frac{1}{n}} \right ),\bar{x}+t_{\alpha}s\left(\sqrt{1+\frac{1}{n}} \right )\right)\\ \\ \\ \left(96- 2.064*1\left(\sqrt{1+\frac{1}{25}} \right ),96+2.064*1\left(\sqrt{1+\frac{1}{25}} \right )\right)\\ \\ \\ (96-2.1049,96+2.1049)\\ \\ \\ (93.8951,98.1049)